plot the supply and demand curve Qd= 3000-10P Qs=-10000 + 10P

1
Expert's answer
2022-04-07T13:08:19-0400

"Q_d=3000-10p,"

"Q_s=-10000+10p,"

"Q_d=Q_s,"

"3000-10p=-10000+10p,"

"13000=20p,"

"p=650,"

"Q_d=Q_s=-3500."